QUESTION ONE

                                                                                           Bank A  

ASSETS

Notes and Coins     69

Balances with Bank of Zambia  30

Balance with Banks and other financial institutions in Zambia   0 

Balance with Banks and other financial institutions abroad       24

Investments in securities   22                  

Loans and advances    57

Fixed Asset 78

Other Assets  17

 

Total Assets  297

 

Deposits  50

Balances due to Banks and other financial institutions in Zambia    88

Balances due to Banks and other financial institutions abroad    14

Other liabilities  58

Bills of Exchange  0

Other borrowed funds (including Subordinated Debt) 71

Shareholders’ equity 35

                                               

Total Liabilities   316

 

Off Balance Sheet items

 

Contingent liabilities                                                                                                                                                         

Assets pledged as collateral other than a house                                                                                                           

Allowance for losses on acceptances and off balance sheet items in other liabilities                                          

Calculate for Banks A: Total Primary Capital, Total Secondary Capital, Eligible Total Capital, current Minimum Total Capital Requirements, and Capital Excess or Deficiency assuming:

a).a foreign majority owned bank                                                                                                                                                             b). a Zambian majority owned bank.              


1
Expert's answer
2021-03-04T07:27:36-0500

Total Primary Capital means, with respect to any Subsidiary, (a) equity plus (b) loan loss reserves minus (c) intangible assets (as set forth on Schedule RC of the call report for such Subsidiary) including mortgage servicing rights.

Total Primary Capital=35+58-0=93

Secondary capital of the bank: term subordinated borrowings and perpetual debt not included in primary capital

Total Secondary Capital=71

Regulatory capital consists of three categories, each governed by a single set of criteria that instruments are required to meet before inclusion in the relevant category.

(1)Common Equity Tier 1 (going-concern capital)

(2)Additional Tier 1 (going-concern capital)

(3)Tier 2 Capital (gone-concern capital)

Eligible Total Capital=93+71=164

The capital ratio is calculated using the definition of regulatory

capital and risk-weighted assets. The total capital ratio must be no lower than 8%. Tier 2

capital is limited to 100% of Tier 1 capital. 

"The capital ratio =\frac{Regulatory capital }{Total Assets}=\frac{164}{297}=0.55"


"Capital Excess =(0.55-0.08)\times297=139.59"